Approximately 19.4% of the U.S population age 5 years and above speaks a language other than English at home. In a large metropolitan area, it was found that 94 residents (age 5 years and above) out of 400 randomly selected residents spoke a language other than English at home. Use the z-test for for a proportion, and test the claim that the proportion for this metropolitan area is higher than the national proportion. Use alpha=0.01 and report the p-value.
Any help with this will be deeply appreciated.

Answer by stanbon(75887)

Approximately 19.4% of the U.S population age 5 years and above speaks a language other than English at home. In a large metropolitan area, it was found that 94 residents (age 5 years and above) out of 400 randomly selected residents spoke a language other than English at home. Use the z-test for for a proportion, and test the claim that the proportion for this metropolitan area is higher than the national proportion. Use alpha=0.01 and report the p-value.
------
Ho: p = 0.194
Ha: p > 0.194 (claim)
-----
sample proportion = p-hat = 94/400 = 0.235
----
test stat:: z(0.235) = (0.235-0.194)/sqrt[0.194*0.806/400] = 2.0737
-----
p-value = P(z > 2.0737) = invNorm(2.0737,100) = 0.0191
----
Conclusion: Since the p-value is greater tha 1% fail to reject Ho.
----
The test results do not support the claim.
======================
Cheers,
Stan H.
======================